Rudy must play 12 commercials his 1-hr radio show Each commercial is either 30 sec or 60 sec long. If the total commercial time during that 1-hr is 10 min how many commercials of each type does Rudy play?
Answer by stanbon(75887)

Rudy must play 12 commercials his 1-hr radio show Each commercial is either 30 sec or 60 sec long. If the total commercial time during that 1-hr is 10 min how many commercials of each type does Rudy play?
--------------------------------------------
Let # of 30 sec commercials be x; time involved is 30x seconds
----------------------------------
# of 60 sec commercials is "12-x" ; time involved is 60(12-x) = 720-60x
----------------------------------
EQUATION:
time + time = 10 min or 600 sec
30x + 720-60x = 600
-30x = -120
x = 4 (# of 30 sec commercials)
12-x = 8 (# of 60 sec commercials)
=========================
Cheers,
Stan H.